1.Sodium hydrogen carbonate, NaHCO3, can be decomposed completely by heating. 2 NaHCO3(s) → Na2CO3(s) + CO2(g) + H2O(g) A sample of impure NaHCO3 with an initial mass of 0.739 g yielded a solid residue (consisting of Na2CO3 and other solids) with a final mass of 0.564 g. Determine the mass percent of NaHCO3 in the sample. 2.In a process for producing acetic acid, oxygen gas is bubbled into acetaldehyde, CH3CHO, containing manganese(II) acetate (catalyst) under pressure at 60°C. 2CH3CHO(l) + O2(g) → 2HC2H3O2(l) In a laboratory test of this reaction, 20.6 g CH3CHO and 13.1 g O2 were put into a reaction vessel. We wish to predict the following: a) How many grams of acetic acid can be produced by this reaction from these amounts of reactants? b) How many grams of the excess reactant remain after the reaction is complete? 3.A sample of phosphonitrylic iodide, PNI2, contains 0.0981 mol of the compound. What is the mass of the sample, in grams?
Answer #1

Loss in man = (0739-0-564) = 0-175 eupmae níおalla we dan 2 male a4aHC@g 6 2 84Y20S 0.175 6 2 0.414 0,474 0.739 64-14 % =